OpenStudy (anonymous):

Find all solutions of the equation in the interval [0, 2pi) cos^3x=cosx

OpenStudy (anonymous):

move cosx to the left side cos^3x - cosx = 0 factor out a cosx cosx(cos^2(x) - 1) = 0 cos^2x - 1 = sin^2(x) and get cosx(sin^2(x)) = 0 so cos(x) = 0 and sin^2(x) = 0 cos(x) is 0 at pi/2 and 3pi/2 and sin^2(x) is 0 when x = 0, and x = pi

OpenStudy (anonymous):

Thanks so much, very helpful and easy to understand. :)

OpenStudy (anonymous):

np glad to help